OpenStudy (anonymous):

Find the leg of each isosceles right triangle when the hypotenuse is of the given measure. Given = 5 square root 6

OpenStudy (studygurl14):

Use the information known about 45-45-90 triangles

OpenStudy (gorv):

isoscelec means both leg will be equal in length

OpenStudy (gorv):



OpenStudy (gorv):

\[(5*\sqrt6)=x^2+x^2\]

OpenStudy (gorv):

2x^2=25*6 2*x^2=150

OpenStudy (gorv):

divide both side by 2 and after that take square root onboth side

OpenStudy (anonymous):

oooooh thanks :)